Skip to main contentSkip to solution

Probability that a man speaks truth is 34\frac{3}{4}. He throws a die and reports that it is a six. The probability that it is actually a six is

Solution

✅ Correct Option: 2

Let E1E_1 = die actually shows a six, E2E_2 = die does not show a six, and AA = man reports that it is a six.

P(truth)=34P(\text{truth}) = \dfrac{3}{4}

P(lies)=1−34=14P(\text{lies}) = 1 - \dfrac{3}{4} = \dfrac{1}{4}

P(E1)=16P(E_1) = \dfrac{1}{6}

P(E2)=56P(E_2) = \dfrac{5}{6}


If the die is a six, the man reports six only when he speaks truth:

P(A∣E1)=34P(A|E_1) = \dfrac{3}{4}

If the die is not a six, the man reports six only when he lies:

P(A∣E2)=14P(A|E_2) = \dfrac{1}{4}


Using Bayes' Theorem:

P(E1∣A)=P(A∣E1)⋅P(E1)P(A∣E1)⋅P(E1)+P(A∣E2)⋅P(E2)P(E_1|A) = \dfrac{P(A|E_1) \cdot P(E_1)}{P(A|E_1) \cdot P(E_1) + P(A|E_2) \cdot P(E_2)}

=34×1634×16+14×56= \dfrac{\frac{3}{4} \times \frac{1}{6}}{\frac{3}{4} \times \frac{1}{6} + \frac{1}{4} \times \frac{5}{6}}

=324324+524= \dfrac{\frac{3}{24}}{\frac{3}{24} + \frac{5}{24}}

=324824= \dfrac{\frac{3}{24}}{\frac{8}{24}}

=38= \dfrac{3}{8}


The probability that it is actually a six is 38\boxed{\dfrac{3}{8}}

Keyboard Shortcuts

  • Left arrow: Previous question
  • Right arrow: Next question
  • S key: Jump to solution
  • Q key: Jump to question